using the numbers 6 2 8 10 5 apply a combination of operations ( addition,subtraction,multiplication,and division) and grouping (PARENTHESES)so that each number is used once,used in the order and the result of the expression is equal to the value of -39
Answer by jim_thompson5910(35256)

This is just one of those problems where you have to try every combination, or hope to get lucky and guess.


It turns out that the answer is
6/2+8-10*5 = -39



Here's my list of combinations (note: I did NOT use parenthesis). You have a scroll a bit to find the combination that's the answer (which is in red).


6+2+8+10+5 =31
6+2+8+10-5 =21
6+2+8+10*5 =66
6+2+8+10/5 =18
6+2+8-10+5 =11
6+2+8-10-5 =1
6+2+8-10*5 =-34
6+2+8-10/5 =14
6+2+8*10+5 =93
6+2+8*10-5 =83
6+2+8*10*5 =408
6+2+8*10/5 =24
6+2+8/10+5 =13.8
6+2+8/10-5 =3.8
6+2+8/10*5 =12
6+2+8/10/5 =8.16
6+2-8+10+5 =15
6+2-8+10-5 =5
6+2-8+10*5 =50
6+2-8+10/5 =2
6+2-8-10+5 =-5
6+2-8-10-5 =-15
6+2-8-10*5 =-50
6+2-8-10/5 =-2
6+2-8*10+5 =-67
6+2-8*10-5 =-77
6+2-8*10*5 =-392
6+2-8*10/5 =-8
6+2-8/10+5 =12.2
6+2-8/10-5 =2.2
6+2-8/10*5 =4
6+2-8/10/5 =7.84
6+2*8+10+5 =37
6+2*8+10-5 =27
6+2*8+10*5 =72
6+2*8+10/5 =24
6+2*8-10+5 =17
6+2*8-10-5 =7
6+2*8-10*5 =-28
6+2*8-10/5 =20
6+2*8*10+5 =171
6+2*8*10-5 =161
6+2*8*10*5 =806
6+2*8*10/5 =38
6+2*8/10+5 =12.6
6+2*8/10-5 =2.6
6+2*8/10*5 =14
6+2*8/10/5 =6.32
6+2/8+10+5 =21.25
6+2/8+10-5 =11.25
6+2/8+10*5 =56.25
6+2/8+10/5 =8.25
6+2/8-10+5 =1.25
6+2/8-10-5 =-8.75
6+2/8-10*5 =-43.75
6+2/8-10/5 =4.25
6+2/8*10+5 =13.5
6+2/8*10-5 =3.5
6+2/8*10*5 =18.5
6+2/8*10/5 =6.5
6+2/8/10+5 =11.025
6+2/8/10-5 =1.025
6+2/8/10*5 =6.125
6+2/8/10/5 =6.005
6-2+8+10+5 =27
6-2+8+10-5 =17
6-2+8+10*5 =62
6-2+8+10/5 =14
6-2+8-10+5 =7
6-2+8-10-5 =-3
6-2+8-10*5 =-38
6-2+8-10/5 =10
6-2+8*10+5 =89
6-2+8*10-5 =79
6-2+8*10*5 =404
6-2+8*10/5 =20
6-2+8/10+5 =9.8
6-2+8/10-5 =-0.2
6-2+8/10*5 =8
6-2+8/10/5 =4.16
6-2-8+10+5 =11
6-2-8+10-5 =1
6-2-8+10*5 =46
6-2-8+10/5 =-2
6-2-8-10+5 =-9
6-2-8-10-5 =-19
6-2-8-10*5 =-54
6-2-8-10/5 =-6
6-2-8*10+5 =-71
6-2-8*10-5 =-81
6-2-8*10*5 =-396
6-2-8*10/5 =-12
6-2-8/10+5 =8.2
6-2-8/10-5 =-1.8
6-2-8/10*5 =0
6-2-8/10/5 =3.84
6-2*8+10+5 =5
6-2*8+10-5 =-5
6-2*8+10*5 =40
6-2*8+10/5 =-8
6-2*8-10+5 =-15
6-2*8-10-5 =-25
6-2*8-10*5 =-60
6-2*8-10/5 =-12
6-2*8*10+5 =-149
6-2*8*10-5 =-159
6-2*8*10*5 =-794
6-2*8*10/5 =-26
6-2*8/10+5 =9.4
6-2*8/10-5 =-0.6
6-2*8/10*5 =-2
6-2*8/10/5 =5.68
6-2/8+10+5 =20.75
6-2/8+10-5 =10.75
6-2/8+10*5 =55.75
6-2/8+10/5 =7.75
6-2/8-10+5 =0.75
6-2/8-10-5 =-9.25
6-2/8-10*5 =-44.25
6-2/8-10/5 =3.75
6-2/8*10+5 =8.5
6-2/8*10-5 =-1.5
6-2/8*10*5 =-6.5
6-2/8*10/5 =5.5
6-2/8/10+5 =10.975
6-2/8/10-5 =0.975
6-2/8/10*5 =5.875
6-2/8/10/5 =5.995
6*2+8+10+5 =35
6*2+8+10-5 =25
6*2+8+10*5 =70
6*2+8+10/5 =22
6*2+8-10+5 =15
6*2+8-10-5 =5
6*2+8-10*5 =-30
6*2+8-10/5 =18
6*2+8*10+5 =97
6*2+8*10-5 =87
6*2+8*10*5 =412
6*2+8*10/5 =28
6*2+8/10+5 =17.8
6*2+8/10-5 =7.8
6*2+8/10*5 =16
6*2+8/10/5 =12.16
6*2-8+10+5 =19
6*2-8+10-5 =9
6*2-8+10*5 =54
6*2-8+10/5 =6
6*2-8-10+5 =-1
6*2-8-10-5 =-11
6*2-8-10*5 =-46
6*2-8-10/5 =2
6*2-8*10+5 =-63
6*2-8*10-5 =-73
6*2-8*10*5 =-388
6*2-8*10/5 =-4
6*2-8/10+5 =16.2
6*2-8/10-5 =6.2
6*2-8/10*5 =8
6*2-8/10/5 =11.84
6*2*8+10+5 =111
6*2*8+10-5 =101
6*2*8+10*5 =146
6*2*8+10/5 =98
6*2*8-10+5 =91
6*2*8-10-5 =81
6*2*8-10*5 =46
6*2*8-10/5 =94
6*2*8*10+5 =965
6*2*8*10-5 =955
6*2*8*10*5 =4800
6*2*8*10/5 =192
6*2*8/10+5 =14.6
6*2*8/10-5 =4.6
6*2*8/10*5 =48
6*2*8/10/5 =1.92
6*2/8+10+5 =16.5
6*2/8+10-5 =6.5
6*2/8+10*5 =51.5
6*2/8+10/5 =3.5
6*2/8-10+5 =-3.5
6*2/8-10-5 =-13.5
6*2/8-10*5 =-48.5
6*2/8-10/5 =-0.5
6*2/8*10+5 =20
6*2/8*10-5 =10
6*2/8*10*5 =75
6*2/8*10/5 =3
6*2/8/10+5 =5.15
6*2/8/10-5 =-4.85
6*2/8/10*5 =0.75
6*2/8/10/5 =0.03
6/2+8+10+5 =26
6/2+8+10-5 =16
6/2+8+10*5 =61
6/2+8+10/5 =13
6/2+8-10+5 =6
6/2+8-10-5 =-4
6/2+8-10*5 =-39
6/2+8-10/5 =9
6/2+8*10+5 =88
6/2+8*10-5 =78
6/2+8*10*5 =403
6/2+8*10/5 =19
6/2+8/10+5 =8.8
6/2+8/10-5 =-1.2
6/2+8/10*5 =7
6/2+8/10/5 =3.16
6/2-8+10+5 =10
6/2-8+10-5 =0
6/2-8+10*5 =45
6/2-8+10/5 =-3
6/2-8-10+5 =-10
6/2-8-10-5 =-20
6/2-8-10*5 =-55
6/2-8-10/5 =-7
6/2-8*10+5 =-72
6/2-8*10-5 =-82
6/2-8*10*5 =-397
6/2-8*10/5 =-13
6/2-8/10+5 =7.2
6/2-8/10-5 =-2.8
6/2-8/10*5 =-1
6/2-8/10/5 =2.84
6/2*8+10+5 =39
6/2*8+10-5 =29
6/2*8+10*5 =74
6/2*8+10/5 =26
6/2*8-10+5 =19
6/2*8-10-5 =9
6/2*8-10*5 =-26
6/2*8-10/5 =22
6/2*8*10+5 =245
6/2*8*10-5 =235
6/2*8*10*5 =1200
6/2*8*10/5 =48
6/2*8/10+5 =7.4
6/2*8/10-5 =-2.6
6/2*8/10*5 =12
6/2*8/10/5 =0.48
6/2/8+10+5 =15.375
6/2/8+10-5 =5.375
6/2/8+10*5 =50.375
6/2/8+10/5 =2.375
6/2/8-10+5 =-4.625
6/2/8-10-5 =-14.625
6/2/8-10*5 =-49.625
6/2/8-10/5 =-1.625
6/2/8*10+5 =8.75
6/2/8*10-5 =-1.25
6/2/8*10*5 =18.75
6/2/8*10/5 =0.75
6/2/8/10+5 =5.0375
6/2/8/10-5 =-4.9625
6/2/8/10*5 =0.1875
6/2/8/10/5 =0.0075